OpenStudy (anonymous):

Newton currently has a balance of $1,716.18 in an account he has held for 29 years. He opened the account with an initial deposit of $784. What is the simple interest rate on the account?

OpenStudy (anonymous):

that doesnt really help seeing as i need to find the rate

OpenStudy (jdoe0001):

hmmmm, s = prt solve for "r"

OpenStudy (anonymous):

so how would i set that up? like I/Pt = Prt/Pt and divide both sides by Pt?

OpenStudy (jdoe0001):

yeap, divide both sides by pt

OpenStudy (jdoe0001):

\(\bf s = prt \implies \cfrac{s}{pt} = r\)

OpenStudy (jdoe0001):

I'm not using the 100 part, thus the amount will come up as decimal

OpenStudy (anonymous):

so it would look like 784*29/784*29? idk how i would set that up :/

OpenStudy (anonymous):

idk i just cant set it up right

OpenStudy (jdoe0001):

well, he started off with $784, thus that's the starting amount, thus is the PRINCIPAL what's the "interest earned"? well, today 29years later, he's got $1716.18 from that original amount the difference between 784 and 1716.18 is 1716.18 - 784 = 932.18

OpenStudy (anonymous):

so 932.18 = and how would i set up the Prt/Pt?

OpenStudy (anonymous):

932.18 = \[\frac{ 784*29 }{784*29 }?\]

OpenStudy (jdoe0001):

\(\bf s = \textit{earned interest} = 932.18\\ p = \textit{principal or original amount deposited}\\ t = years\\\qquad \\ s = prt \implies \cfrac{s}{pt} = r \implies \cfrac{932.18}{784\times 29} = r \)

OpenStudy (anonymous):

ok i got 34.48

OpenStudy (jdoe0001):

34.48? ... recheck your calculation :|

OpenStudy (anonymous):

i typed everything right im pretty sure

OpenStudy (jdoe0001):

since we're not using the "100" figure as the one in the picture, the amount for the rate will show as a decimal number, less than 1

OpenStudy (anonymous):

i typed this into my calculator 932.18/784*29 is that not right?

OpenStudy (jdoe0001):

keep in mind, PEMDAS :)

OpenStudy (jdoe0001):

932.18/784 then * 29, yes is 34.48

OpenStudy (anonymous):

then how would i type that? :/ im sorry i dont understand

OpenStudy (jdoe0001):

for a calculator, you'd need to group it => 932.18/(784*29) so the denominator gets evaluated, before any division takes place

OpenStudy (anonymous):

oh alright so not i got 0.041 so it would be 4.1%?

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

:^)

OpenStudy (anonymous):

thanks friend

OpenStudy (jdoe0001):

yw